INAF queue observing, 2013 Nov02-Nov03

Observer: M. Fumana, R. Speziali
Telescope Operator: G. Becchetti
Support Astronomer: H. Khun (on call)
Instrument: MODS (LUCI)

Summary

At the sunset thin cirrus at the horizon.
Cloud and wind come during the night.
Mismatch ID in the Lardo masks. Try to observe Israel program. Had to stop due to bad weather conditions

Rain spots after 5:30.
Snow in the second part of the night.

Data Summary

Project_RUN Objects Filter ExpTime Notes

Details

00:30 The TO opens the dome.

01:09 We send the acquisition, even if we are not yet collimated. We want test the grism tilt in LUCI

01:34 executeLUCIScript.sh test_campoA_K_2.14.spec
  • Grism tilt fails

01:36 executeLUCIScript.sh test_campoA_K_2.15.spec
  • Grism tilt fails

01:39 test_16
  • Grism ok using central wavelength 1.6. But 1.6 is bigger than 1.5 of the previous mask. Abort.

01:43 Switch to MODS

02:10 Slew to Lardo
  • acqMODS M15RGBf1.acq
  • Problem with the mask ID. The MODS slot ID doesn't match the scripts ID

02:42 Slew to Israel 28_A4
  • acqMODS catsj204734_ls.acq

02:50 We increase the exposure time in the acquisition image

03:03 No object found in the image. We try to change target

03:06 Slew to Israel 28_A5: acqMODS catsj215447_ls.acq

03:21 execMODS catsj215447_ls.obs
  • seeing ~1.2"
  • seeing bad ~1.8"-2"
  • clouds are coming
  • seeing >2.5"
  • wind is raising
  • script aborted after 3 exposures
  • strong forces in the mirror due to the worst seeing

05:12 Close the dome.

-- MarcoFumana - 03 Nov 2013
Topic revision: r2 - 05 Nov 2013, RobertoSpeziali
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback